OpenStudy (anonymous):

differentiate dy/dx = xsqrt(y)cos^2(sqrt(y))

myininaya (myininaya):

So you want to find y'' given y' so we have \[y'=x \sqrt{y} \cos^2(\sqrt{y})\]

OpenStudy (anonymous):

no i think you need to find y

OpenStudy (anonymous):

the directions said solve for y

myininaya (myininaya):

so we want to solve the differential equation not differentiate ?

myininaya (myininaya):

so this is a separation of variables problem

OpenStudy (zarkon):

it is a separable diffeq...not too bad

myininaya (myininaya):

\[\frac{1}{\sqrt{y} \cos^2(\sqrt{y})}dy=x dx\]

myininaya (myininaya):

integrate both sides

myininaya (myininaya):

don't forget to put +C on of the sides

myininaya (myininaya):

one*

OpenStudy (anonymous):

how would you integrwte the left side though? thats what confuses me

myininaya (myininaya):

that one side is not that bad with a substitution u=sqrt{y} and write 1/cos^2 as sec^2

myininaya (myininaya):

\[u=\sqrt{y} => du=\frac{1}{2 \sqrt{y}} dy => 2 du=\frac{dy}{\sqrt{y}}\]

myininaya (myininaya):

\[\int\limits_{}^{} 2 \sec^2(u) du\]

myininaya (myininaya):

\[2 \tan(u)+C=2 \tan(\sqrt{y})+C\]

myininaya (myininaya):

did i lose you?

OpenStudy (anonymous):

no thank you so much that really helped

myininaya (myininaya):

ok you still have to do the other side don't forget

OpenStudy (anonymous):

wait so how would you format your final answer?

myininaya (myininaya):

\[2 \tan(\sqrt{y})+C=\frac{x^2}{2}\]

myininaya (myininaya):

I would leave it like this

myininaya (myininaya):

I think Some teachers like you to solve it for y I think

OpenStudy (anonymous):

yeah that would be a pain but thanks for all your help

OpenStudy (zarkon):

I would leave it like that too.

myininaya (myininaya):

Zarkon has a phd so listen to him.

OpenStudy (anonymous):

okay that saves a lot of work

OpenStudy (zarkon):

since the tangent is periodic when you take the inverse tangent you might not reclaim the y value you wanted.

myininaya (myininaya):

right moving stuff around sometimes messes with domain right?

OpenStudy (zarkon):

yes...best to leave it as it is

myininaya (myininaya):

just like you aren't suppose to solve this x=x^2 by dividing x on both sides